OpenStudy (anonymous):

the following info is known about a cubic polynomial function: p(-1)=0, p(3)= 0, p(6)=0 and p(5)=48. Find the equation of the function in factored form.

OpenStudy (swissgirl):

Since at x=1,3and 6 y=0 Our roots will be at 1,-3 and -6 p(x)=a(x+1)(x-3)(x-6) Solve for a: 48=a(5+1)(5-3)(5-6) Once you solve for a .... Voila you have your equation

OpenStudy (anonymous):

I get it. Thank you soooo much! Can i ask you another question?

OpenStudy (swissgirl):

yes

OpenStudy (anonymous):

What are the potential integral zeros of the polynomial p(x)=6x^3-2x^2-x+4?

OpenStudy (swissgirl):

Not sure what potential integral zeros mean

OpenStudy (anonymous):

me either ;(

OpenStudy (swissgirl):

@bibby

OpenStudy (bibby):

idk google it nerd

OpenStudy (swissgirl):

I did a$$hole .

OpenStudy (bibby):

bby pls

OpenStudy (anonymous):

i think it's something about the leading coefficient and constant term.

OpenStudy (swissgirl):

k gonna google it

OpenStudy (swissgirl):

Ohh found it :)

OpenStudy (anonymous):

really?!

OpenStudy (swissgirl):

a= factors of constant term which is 4 a=1,2,4 b= factors of leading coefficient which is 6 b=1,2,3,6

OpenStudy (anonymous):

what do i do with them?

OpenStudy (swissgirl):

Potential Integral Zeroes are +-a/b

OpenStudy (anonymous):

So the potential integral zeroes would be 1, 1/2, 1/3, 1/6,2,2/3,4,4/3

OpenStudy (swissgirl):

1,2,4 1/2,1, 4/2 1/3,2/3,4/3 1/6,2/6,4/6 We can simplify it to ... 1,2,4, 1/2,1/3,2/3,4/3,1/6

OpenStudy (anonymous):

thank you soooooo much for the help! :)

OpenStudy (swissgirl):

np

OpenStudy (anonymous):

can I ask you another question if you don't mind? I have a test tomorrow.

OpenStudy (swissgirl):

ya sure

OpenStudy (anonymous):

i posted the question.

OpenStudy (aum):

"potential integral zeros of the polynomial" INTEGRAL zeros means zeros that are INTEGERS. 6x^3-2x^2-x+4 For potential integer zeros you need to consider only the factors of the constant term. Here the constant term is 4. And the factors of 4 are: 1, 2 and 4. The potential integral zeros are: 1, 2 and 4. There is another theorem called the rational roots theorem where you consider the factors of the constant term divided by the factors of the leading coefficient. But that is not what they are asking here.